# POGEMA: A Benchmark Platform for Cooperative Multi-Agent Navigation

### v1

@@ Abstract
Multi-agent reinforcement learning (MARL) has recently excelled in solving challenging cooperative and competitive multi-agent problems in various environments with, mostly, few agents and full observability. Moreover, a range of crucial robotics-related tasks, such as multi-robot navigation and obstacle avoidance, that have been conventionally approached with the classical non-learnable methods (e.g., heuristic search) is currently suggested to be solved by the learning-based or hybrid methods. Still, in this domain, it is hard, not to say impossible, to conduct a fair comparison between classical, learning-based, and hybrid approaches due to the lack of a unified framework that supports both learning and evaluation. To this end, we introduce POGEMA, a set of comprehensive tools that includes a fast environment for learning, a generator of problem instances, the collection of pre-defined ones, a visualization toolkit, and a benchmarking tool that allows automated evaluation. We introduce and specify an evaluation protocol defining a range of domain-related metrics computed on the basics of the primary evaluation indicators (such as success rate and path length), allowing a fair multi-fold comparison. The results of such a comparison, which involves a variety of state-of-the-art MARL, search-based, and hybrid methods, are presented.
@@
Figure 1: (a) Example of the multi-robot navigation problem considered in POGEMA: each robot must reach its goal, denoted by a flag of the same color. (b) Observation tensor of the red agent. (c) Results of the evaluation of several MARL, hybrid, and search-based solvers on the proposed POGEMA benchmark.
@@ 1 Introduction
Multi-agent reinforcement learning (MARL) has gained an increasing attention recently and significant progress in this field has been achieved [6, 32, 62]. MARL methods have been demonstrated to generate well-performing agents’ policies in strategic games [2, 63], sport simulators [49, 66], multi-component robot control [59], city traffic control [20], and autonomous driving [70]. Currently, several ways to formulate and solve MARL problems exist, based on what information is available to the agents and what type of communication is allowed in the environment [67]. Due to the increased interest in robotic applications, decentralized cooperative learning with minimizing communication between agents has recently attracted a specific attention [45, 68]. Decentralized learning naturally suits the partial observability of the environment in which the robots usually operate. Reducing the information transmitted through the communication channels between the agents increases their degree of autonomy.
The main challenges in solving MARL problems are the non-stationarity of the multi-agent environment, the need to explicitly predict the behavior of the other agents to implement cooperative behavior, high dimensionality of the action space, which grows exponentially with the number of agents, and the sample inefficiency of existing approaches. The existing MARL including model-based and hybrid learnable methods [12, 26] exhibit faster and more stable learning in SMAC-type environments [13] with vector observations and full observability. Currently, the best results are shown by the discrete explicit world models, that use Monte Carlo tree search for planning with various heuristics to reduce the search space [18, 26].
However, in numerous practically inspired applications, like in mobile robot navigation, agents’ observations are typically high-dimensional (e.g. stacked occupancy grid matrices or image-based observations as compared to 32-dim vectors in SMAC [13]) and only partially describe the state of the environment, including the other agents [17, 14]. This makes the problem specifically challenging, especially in the environments where a large number of agents are involved. For example, it is not uncommon in robotics to consider settings where up to hundreds of agents are acting (moving) simultaneously in the shared workspace as opposed to 2–10 agents in conventional MARL environments such as SMAC [13] or Google Research Football [21]. Learning to act in such crowded, observation-rich and partially-observable environments is a notable challenge to existing MARL methods.
Conventionally, the problem of multi-robot cooperative navigation (which is very important due to its applications in modern automated warehouses and fulfillment centers [11]) is framed as a search problem over a discretized search space, composed of robots-locations tuples. All robots are assumed to be confined to a graph, typically – a 4-connected grid [38], and at each time step a robot can either move following a graph’s edge or stay at the current vertex. This problem setting is known as (Classical) Multi-agent Pathfinding problem [50]. Even in such simplified setting (discretized space, discretized time, uniform-duration actions etc.) obtaining a set of individual plans (one for each robot) that are mutually-conflict-free (i.e. no vertex or edge is occupied by disctinct agents at the same time step) and minimize a common objective such as, for example, the arrival time of the last agent (known as the makespan in the literature) is NP-Hard [65]. Moreover if the underlying graph is directed even obtaining a valid solution is HP-Hard as well [31].
@@
To this end the focus of the multi-agent pathfinding community is recently being shifted towards exploring of how state-of-the-art machine learning techniques, especially reinforcement learning and imitation learning, can be leveraged to increase the efficiency of traditional solvers. Methods like [48, 46, 47, 29, 61, 42, 58, 27, 10] are all hybrid solvers that rely on both widespread search-based techniques and learnable components as well. They all are developed using different frameworks, environments and datasets and are evaluated accordingly, i.e. in the absence of the unifying evaluation framework, consisting of the (automated) evaluation tool, protocol (that defines common performance indicators) and the dataset of the problem instances. Moreover, currently most of the pure MARL methods, i.e. the ones that do not involve search-based modules, such as QMIX [37], MAMBA [12], MAPPO [64] etc., are mostly not included in comparison. The main reason is that to train MARL policies a fast environment is needed, which is suited to cooperative multi-agent navigation.
@@
  • a fast and flexible environment for learning and planning supporting several variants of the multi-robot navigation problem,
  • a generator of problem instances for multi-task and generalization testing,
  • a visualization toolkit to create plots for debugging and performance information and to make high-quality animations,
  • a benchmarking tool that allows automated evaluation of both learnable, planning, and hybrid approaches.
Moreover, we introduce and specify an evaluation protocol defining a range of domain-related metrics computed on the basics of the primary evaluation indicators (such as success rate and path length), allowing a fair multi-fold comparison of learnable and classical methods. The results of such a comparison, which involves a range of the state-of-the-art MARL, search-based, and hybrid methods, are presented. The source code for POGEMA Benchmark is available at the following links [^1] https://github.com/AIRI-Institute/pogema-benchmark [^2] https://github.com/AIRI-Institute/pogema-toolbox [^3] https://github.com/AIRI-Institute/pogema. Additionally, the POGEMA environment and pogema-toolbox can be installed via PyPI using pip.
@@ 2 Related Work
Currently, a huge variety of MARL environments exists that are inspired by various practical applications and encompass a broad spectrum of nuances in problem formulations. Notably, they include a diverse array of computer games [41, 13, 39, 7, 51, 19, 5, 3, 21]. Additionally, they address complex social dilemmas [1] including public goods games, resource allocation problems [35], and multi-agent coordination challenges. Some are practically inspired, showcasing tasks such as competitive object tracking [34], infrastructure management and planning [23], and automated scheduling of trains [30]. Beyond these, the environments simulate intricate, interactive systems such as traffic management and autonomous vehicle coordination [57], multi-agent control tasks [39, 36], and warehouse management [16]. Each scenario is designed to challenge and analyze the collaborative and competitive dynamics that emerge among agents in varied and complex contexts. We summarize the most wide-spread MARL environments in Table 1. A detailed description of each column is presented below.
Table 1: Comparison of different multi-agent reinforcement learning environments
@@
Flatland [30] link Coop
GoBigger [69] link Mixed/Coop
Google Research Football [21] link Mixed
Griddly [4] link Mixed
Hide-and-Seek [3] link Comp
IMP-MARL [23] link Coop
Jumanji (XLA) [5] link Mixed
LBF [35] link Coop
MAMuJoCo [36] link Coop
MATE [34] link Coop
MeltingPot [1] link Mixed/Coop
Minecraft MALMO [19] link Mixed
MPE [28] link Mixed
MPE (XLA) [39] link Mixed
Multi-agent Brax (XLA) [39] link Coop
Multi-Car Racing [44] link Comp
Neural MMO [51] link Comp
Nocturne [57] link Mixed
Overcooked [7] link Coop
Overcooked (XLA) [39] link Coop
RWARE [35] link Coop
SISL [16] link Coop
SMAC [41] link Mixed/Coop
SMAC v2 [13] link Mixed/Coop
SMAX (XLA) [39] link Mixed/Coop
@@
Python based
Most of the current multi-agent environments are implemented in Python, however, some of them include bindings of the other program languages or external dependencies that might complicate their usage. Pure Python implementations generally ensure ease of modification and customization, allowing researchers to readily adapt and extend the environments.
Hardware-agnostic setup
means that the environment doesn’t require setup for any specific type of training or inferencing hardware.
@@
To improve the ability of RL agents to solve problem instances that are not the same that were used for training (the so-called ability to generalize) procedural generation of the problem instances is commonly used. I.e. the environment does not rely on a predefined set of training instances but rather procedurally generates them to prevent overfitting. As highlighted in the Procgen paper [9], this approach ensures that agents develop robust strategies capable of adapting to novel and diverse situations. Moreover, in multi-agent settings, agents must be able to handle and adapt to a variety of unforeseen agent behaviors and strategies, ensuring robustness and flexibility in dynamic environments [1].
@@
Tests and CI
means that the environment is set up for development with continuous integration and is covered with tests, which are essential for collaborative open-source development.
PyPI listed
means that the environment library is listed on PyPI [^4] https://pypi.org, making it easy to install and integrate into your projects with a simple pip install command.
Scalability to >1000 Agents
stands for the environment’s ability to handle more than 1000 simultaneously acting agents, ensuring robust performance and flexibility for large-scale multi-agent simulations.
Induced behaviour
Multi-agent behaviour can be induced by modifying the reward function. Competitive (Comp) behaviour occurs when all agents’ rewards sum to a fixed constant at each time step. As an example, minimax games, where rewards sum up to zero, are competitive games. Cooperative/collaborative (Coop) behaviour happens when the agents share a reward function – e.g. when both agents have the same goal in the environment and get rewarded only by its completion. Mixed behaviour (Mixed) does not constrain the reward function in any way and can be described as a combination of cooperative and competitive behaviours. One of the most famous examples of these games is iterated prisoner’s dilemma.
As we aim to create a lightweight and easy-to-configure multi-agent environment for reinforcement learning and pathfinding tasks, we consider the following factors essential. First and foremost, our environment is fully compatible with the native Python API: we target pure Python builds independent of hardware-specific software with a minimal number of external dependencies. Moreover, we underline the importance of constant extension and flexibility of the environment. Thus, we prioritize testing and continuous integration as cornerstones of the environment, as well as trouble-free modification of the transition dynamics. Secondly, we highlight that our environment targets generalization and may utilize procedural generation. Last but not least, we target high computational throughput (i.e., the number of environment steps per second) and robustness to an extremely large number of agents (i.e., the environment remains performant under high loads).
There are many environments inducing various types of multi-agent behaviors via different reward structures. Unfortunately, many of them require extensive Python support and rely on APIs of different programming languages (e.g., Lua, C++) for lower latency or depend on hardware-specific libraries such as XLA. Furthermore, many environments do not support generalization and lack procedural generation, especially in multi-agent cases. Additionally, customization of certain environments might be considered an issue without reverse engineering them. That’s why we emphasize the superiority of the proposed benchmark.
@@
The evaluation protocols in these environments typically feature learning curves that highlight the performance of each algorithm under specific scenarios. For SMAC, these scenarios involve games against predefined bots with specific units on both sides. In MAMuJoCo, the standard tasks involve agents controlling different sets of joints, while in GRF, the scenarios are games against predefined policies from Football Academy scenarios. Proper evaluation of MARL approaches is a serious concern. For SMAC, it’s highlighted in the paper [15], which proposes a unified evaluation protocol for this benchmark. This protocol includes default evaluation parameters, performance metrics, uncertainty quantification, and a results reporting scheme.
@@ 3 POGEMA
POGEMA, which comes from Partially-Observable Grid Environment for Multible Agents, is an umbrella name for a collection of versatile and flexible tools aimed at developing, debugging and evaluating different methods and policies tailored to solve several types of multi-agent navigation tasks.
@@ 3.1 POGEMA Environment
POGEMA [^5] https://pypi.org/project/pogema environment is a core of POGEMA suite. It implements the basic mechanics of agents’ interaction with the world. The environment can be installed using the Python Package Index (PyPI). The environemnt is open-sourced and available at github [^6] https://github.com/AIRI-Institute/pogema under MIT license. POGEMA provides integration with existing RL frameworks: PettingZoo [54], PyMARL [40], and Gymnasium [56].
@@
POGEMA supports two generic types of multi-agent navigation problems. In the first variant, dubbed MAPF (from Multi-agent Pathfinding), each agent is provided with the unique goal location and has to reach it avoiding collisions with the other agents and static obstacles. For MAPF problem setting POGEMA supports both stay-at-target behavior (when the episode successfully ends only if all the agents are at their targets) and disappear-at-target (when the agent is removed from the environment after it first reaches its goal). The second variant is a lifelong version of multi-agent navigation and is dubbed accordingly – LMAPF. Here each agent upon reaching a goal is immediately assigned another one (not known to the agent beforehand). Thus the agents are constantly moving trough in the environment until episode ends.
@@
At each timestep each agent in POGEMA receives an individual ego-centric observation represented as a tensor – see Fig. 1. The latter is composed of the following $(2R+1)\times(2R+1)$ binary matrices, where $R$ is the observation radius set by the user:
@@
The suggested observation, which is, indeed, minimalist and simplistic, can be modified by the user using wrapper mechanisms. For example, it is not uncommon in the MAPF literature to augment the observation with additional matrices encoding the agent’s path-to-goal (constructed by some global pathfinding routine) [46] or other variants of global guidance [29].

### v2

# POGEMA: A Benchmark Platform for Cooperative Multi-Agent Pathfinding

@@ Abstract
Multi-agent reinforcement learning (MARL) has recently excelled in solving challenging cooperative and competitive multi-agent problems in various environments, typically involving a small number of agents and full observability. Moreover, a range of crucial robotics-related tasks, such as multi-robot pathfinding, which have traditionally been approached with classical non-learnable methods (e.g., heuristic search), are now being suggested for solution using learning-based or hybrid methods. However, in this domain, it remains difficult, if not impossible, to conduct a fair comparison between classical, learning-based, and hybrid approaches due to the lack of a unified framework that supports both learning and evaluation. To address this, we introduce POGEMA, a comprehensive set of tools that includes a fast environment for learning, a problem instance generator, a collection of predefined problem instances, a visualization toolkit, and a benchmarking tool for automated evaluation. We also introduce and define an evaluation protocol that specifies a range of domain-related metrics, computed based on primary evaluation indicators (such as success rate and path length), enabling a fair multi-fold comparison. The results of this comparison, which involves a variety of state-of-the-art MARL, search-based, and hybrid methods, are presented.
@@
Figure 1: (a) Example of the multi-agent pathfinding problem considered in POGEMA: each agent must reach its goal, denoted by a flag of the same color. (b) Observation tensor of the red agent. (c) Evaluation results of MARL, hybrid, and search-based solvers on POGEMA benchmark.
@@ 1 Introduction
Multi-agent reinforcement learning (MARL) has gained increasing attention recently, with significant progress achieved in the field (Canese et al., 2021; Nguyen et al., 2020; Wong et al., 2023). MARL methods have demonstrated the ability to generate well-performing agents’ policies in strategic games (Arulkumaran et al., 2019; Ye et al., 2020), sports simulators (Song et al., 2023; Zang et al., 2024), multi-component robot control (Wang et al., 2024), city traffic control (Kolat et al., 2023), and autonomous driving (Zhou et al., 2020). Several approaches currently exist for formulating and solving MARL problems, depending on the information available to agents and the type of communication allowed in the environment (Zhang et al., 2021). With the growing interest in robotic applications, decentralized cooperative learning that minimizes communication between agents has recently attracted particular attention (Singh et al., 2022; Zhang et al., 2020).
A prominent example of an important and practically-inspired problem that can benefit from this type of learning is the so-called multi-agent pathfinding (MAPF) (Stern et al., 2019). In this problem a group of (homogeneous) agents is confined to a graph of locations and at each time step an agent can either wait at the current vertex or move to an adjacent one. Each agent is assigned a goal and the task is to make the agents reach their goals as fast as possible (i.e. using fewer actions) in a safe way, i.e. avoiding the inter-agent collisions as well as collisions with the static obstacles. Numerous practically important applications mimic the discrete nature of MAPF and actively exploit MAPF solutions in real world. A major example is automated warehouses (Dekhne et al., 2019) where robots move synchronously utilizing atomic moves. Furthermore, many works, e.g., (Hönig et al., 2016; Ma et al., 2019; Okumura et al., 2022), describe how MAPF solutions can be transferred to real robots that are subject to kinodynamic constraints, inaccurate execution, and other physically inspired complications. Thus, the MAPF problem serves as a highly useful abstraction that distills the core challenge of any multi-agent navigation problem: coordinating actions between agents to minimize the risk of collisions while optimizing a given cost objective.
Conventional MAPF solvers rely on such algorithmic techniques as systematic (heuristic) search: A*+ID+OD (Standley, 2010), CBS (Sharon et al., 2015), M* (Wagner & Choset, 2011), ICTS (Sharon et al., 2013); or reduce the MAPF problem to the other established computational problem like boolean satisfiability (Surynek et al., 2016) or integer linear programming; or leverage dedicated rule-based techniques (de Wilde et al., 2013). Meanwhile, learning-based MAPF solvers have recently been receiving increased attention, such as (Sartoretti et al., 2019; Ma et al., 2021; Damani et al., 2021; Skrynnik et al., 2024a), to name a few. A key advantage of such solvers is their decentralized nature, allowing each agent to act independently, which can significantly reduce costs compared to classical MAPF solvers requiring centralized control.
From a learning perspective, MAPF is a highly specific problem that differs significantly from well-known MARL problems such as the StarCraft Multi-agent Challenge (SMAC) (Ellis et al., 2024) or Google Research Football (Kurach et al., 2020). In MAPF, generalization to different types of maps and numbers of agents is essential, as real-world MAPF solvers must handle varying map topologies and agent populations. Furthermore, the number of agents in MAPF is often very large – not just dozens (as in SMAC) but hundreds or even thousands of agents are moving simultaneously in the environment. Unsurprisingly, the vast majority of state-of-the-art learnable MAPF solvers, such as (Skrynnik et al., 2023; 2024a; 2024b; Ma et al., 2021; Wang et al., 2023; Sartoretti et al., 2019; Wang et al., 2020a; Liu et al., 2020; Damani et al., 2021), are the hybrid solvers that rely on both traditional search-based techniques and learnable components.
@@
These solvers are developed using different frameworks, environments and datasets and are evaluated accordingly, i.e. in the absence of the unifying evaluation framework, consisting of the evaluation tool, protocol (that defines common performance indicators) and the dataset of the problem instances. Moreover, currently most of the pure MARL methods, i.e. the ones that do not involve search-based modules, such as QMIX (Rashid et al., 2020), MAMBA (Egorov & Shpilman, 2022), MAPPO (Yu et al., 2022) etc., are mostly not included in comparison. This exclusion is likely due to the lack of a unified benchmark that is compatible with or integrated with existing MARL learning frameworks.
@@
  • a fast and flexible environment supporting different multi-robot pathfinding problems, coupled with a generator of diverse problem instances to facilitate multi-task learning and generalization testing,
  • a visualization toolkit enabling the creation of high-quality vector-based plots and animations for enhanced analysis and presentation,
  • a benchmarking tool for automated, parallel evaluation of learning-based, planning-based, and hybrid approaches, streamlining comparison across methodologies,
  • a standardized evaluation protocol offering domain-specific metrics derived from primary performance indicators, ensuring robust and fair comparisons between methods.
@@ 2 Related Work
Currently, a huge variety of MARL environments exists that are inspired by various practical applications and encompass a broad spectrum of nuances in problem formulations. Notably, they include a diverse array of computer games (Samvelyan et al., 2019b; Ellis et al., 2024; Rutherford et al., 2023; Carroll et al., 2019; Suarez et al., 2024; Johnson et al., 2016; Bonnet et al., 2023; Baker et al., 2020; Kurach et al., 2020). Additionally, they address complex social dilemmas (Agapiou et al., 2022) including public goods games, resource allocation problems (Papoudakis et al., 2021), and multi-agent coordination challenges. Some are practically inspired, showcasing tasks such as competitive object tracking (Pan et al., 2022), infrastructure management and planning (Leroy et al., 2024), and automated scheduling of trains (Mohanty et al., 2020). Beyond these, the environments simulate intricate, interactive systems such as traffic management and autonomous vehicle coordination (Vinitsky et al., 2022), multi-agent control tasks (Rutherford et al., 2023; Peng et al., 2021), and warehouse management (Gupta et al., 2017). Each scenario is designed to challenge and analyze the collaborative and competitive dynamics that emerge among agents in varied and complex contexts. We summarize the most wide-spread MARL environments in Table 1. A detailed description of each column is presented below.
@@
Flatland (Mohanty et al., 2020) link Coop
Gigastep (Lechner et al., 2024) link Mixed
GoBigger (Zhang et al., 2023) link Mixed/Coop
Google Research Football (Kurach et al., 2020) link Mixed
Griddly (Bamford, 2021) link Mixed
Hide-and-Seek (Baker et al., 2020) link Comp
IMP-MARL (Leroy et al., 2024) link Coop
Jumanji (XLA) (Bonnet et al., 2023) link Mixed
LBF (Papoudakis et al., 2021) link Coop
MAMuJoCo (Peng et al., 2021) link Coop
Magent (Zheng et al., 2018) link Mixed
MATE (Pan et al., 2022) link Coop
MeltingPot (Agapiou et al., 2022) link Mixed/Coop
MALMO (Johnson et al., 2016) link Mixed
MPE (Lowe et al., 2017) link Mixed
MPE (XLA) (Rutherford et al., 2023) link Mixed
Multi-agent Brax (XLA) (Rutherford et al., 2023) link Coop
Multi-Car Racing (Schwarting et al., 2021) link Comp
Neural MMO (Suarez et al., 2024) link Comp
Nocturne (Vinitsky et al., 2022) link Mixed
Overcooked (Carroll et al., 2019) link Coop
Overcooked (XLA) (Rutherford et al., 2023) link Coop
RWARE (Papoudakis et al., 2021) link Coop
SISL (Gupta et al., 2017) link Coop
SMAC (Samvelyan et al., 2019b) link Mixed/Coop
SMAC v2 (Ellis et al., 2024) link Mixed/Coop
SMAX (XLA) (Rutherford et al., 2023) link Mixed/Coop
@@
Table 1: Comparison of different multi-agent reinforcement learning environments.
@@
Environment Setup
Python based means the environment is implemented in Python, unlike many other multi-agent environments that include bindings for other languages or external dependencies, which can complicate usage. Pure Python implementations ensure ease of modification and customization, allowing researchers to easily adapt and extend the environments. Hardware-agnostic setup means the environment doesn’t require any specific type of hardware for training or inference, offering flexibility across different systems.
@@
To improve the ability of RL agents to solve problem instances that are not the same that were used for training (the so-called ability to generalize) procedural generation of the problem instances is commonly used. I.e. the environment does not rely on a predefined set of training instances but rather procedurally generates them to prevent overfitting. As highlighted in the Procgen paper (Cobbe et al., 2020), this approach ensures that agents develop robust strategies capable of adapting to novel and diverse situations. Moreover, in multi-agent settings, agents must be able to handle and adapt to a variety of unforeseen agent behaviors and strategies, ensuring robustness and flexibility in dynamic environments (Agapiou et al., 2022).
@@
Tests and CI means the environment is set up for development with continuous integration and is covered with tests, which are essential for collaborative open-source development. PyPI listed indicates that the environment library is listed on PyPI [^1] https://pypi.org, making it easy to install and integrate into projects with a simple pip install command. Scalability to >1000 Agents refers to the environment’s ability to handle over 1000 simultaneously acting agents, ensuring robust performance and flexibility for large-scale multi-agent simulations.
Induced behavior
Multi-agent behavior can be influenced by modifying the reward function (Shoham & Leyton-Brown, 2008). Competitive (Comp) behavior arises when a joint strategy benefits one player but disadvantages others. In a two-player game, this corresponds to a Pareto-efficient outcome. Minimax games, where agents’ rewards sum to zero, are classic examples of competitive games. Cooperative (Coop) behavior (Du et al., 2023; Shoham & Leyton-Brown, 2008) occurs when agents share a unified reward function or pursue the same goal, rewarded only by its completion. Social dilemmas are a key example of cooperation. Mixed behavior (Littman, 1994) doesn’t limit the agents’ objectives or interactions, blending cooperative and competitive behaviors. A well-known example is the iterated prisoner’s dilemma.
As we aim to create a lightweight and easy-to-configure multi-agent environment for reinforcement learning and pathfinding tasks, we consider the following factors essential. First and foremost, our environment is fully compatible with the native Python API: we target pure Python builds independent of hardware-specific software with a minimal number of external dependencies. Moreover, we underline the importance of constant extension and flexibility of the environment. Thus, we prioritize testing and continuous integration as cornerstones of the environment, as well as trouble-free modification of the transition dynamics. Secondly, we highlight that our environment targets generalization and may utilize procedural generation. Last but not least, we target high computational throughput (i.e., the number of environment steps per second) and robustness to an extremely large number of agents (i.e., the environment remains efficient under high loads). The detailed overview of the mentioned related papers is provided in Appendix F.
@@
The evaluation protocols in these environments typically feature learning curves that highlight the performance of each algorithm under specific scenarios. For SMAC, these scenarios involve games against predefined bots with specific units on both sides. In MAMuJoCo, the standard tasks involve agents controlling different sets of joints, while in GRF, the scenarios are games against predefined policies from Football Academy scenarios. Proper evaluation of MARL approaches is a serious concern. For SMAC, it’s highlighted in the paper (Gorsane et al., 2022), which proposes a unified evaluation protocol for this benchmark. This protocol includes default evaluation parameters, performance metrics, uncertainty quantification, and a results reporting scheme.
@@ 3 POGEMA
POGEMA, which comes from Partially-Observable Grid Environment for Multiple Agents, is an umbrella name for a collection of versatile and flexible tools aimed at developing, debugging and evaluating different methods and policies tailored to solve several types of multi-agent pathfinding tasks. The source code is available at: POGEMA Benchmark [^2] https://github.com/Cognitive-AI-Systems/pogema-benchmark, POGEMA Toolbox [^3] https://github.com/Cognitive-AI-Systems/pogema-toolbox and POGEMA Environment [^4] https://github.com/Cognitive-AI-Systems/pogema.
@@ 3.1 POGEMA Environment
POGEMA  [^5] https://pypi.org/project/pogema environment is a core of POGEMA suite. It implements the basic mechanics of agents’ interaction with the world. The environemnt is open-sourced under MIT license. POGEMA provides integration with existing RL frameworks: PettingZoo (Terry et al., 2021), PyMARL (Samvelyan et al., 2019a), and Gymnasium (Towers et al., 2023).
@@
POGEMA supports two generic types of multi-agent navigation problems. In the first variant, dubbed MAPF (from Multi-agent Pathfinding), each agent is provided with the unique goal location and has to reach it avoiding collisions with the other agents and static obstacles. For MAPF problem setting POGEMA supports both stay-at-target behavior (when the episode successfully ends only if all the agents are at their targets) and disappear-at-target (when the agent is removed from the environment after it first reaches its goal). The second variant is a lifelong version of multi-agent pathfinding and is dubbed accordingly – LMAPF. Here each agent upon reaching a goal is immediately assigned another one (not known to the agent beforehand). Thus, the agents are constantly moving through in the environment until episode ends.
@@
At each timestep each agent in POGEMA receives an individual ego-centric observation represented as a tensor – see Figure 1. The latter is composed of the following $(2R+1)\times(2R+1)$ binary matrices, where $R$ is the observation radius set by the user:
@@
The suggested observation, which is, indeed, minimalist and simplistic, can be modified by the user using wrapper mechanisms. For example, it is not uncommon in the MAPF literature to augment the observation with additional matrices encoding the agent’s path-to-goal (constructed by some global pathfinding routine) (Skrynnik et al., 2024a) or other variants of global guidance (Ma et al., 2021).